x^2(y-1)uxx-x(y^2-1)uxy+y(y-1)uyy+(xy)ux-uy =0
Reduce into canonical form and solve it
1
Expert's answer
2020-12-15T02:41:06-0500

Transforming the PDE into a canonical form.


Comparing the given partial differential equation with


"A(x,y)u_{xx}+B(x,y)u_{xy}+C(x,y)u_{yy}+D(x,y)u_x+E(x,y)u_y+F(x,y)u+G(x,y)= 0,"


We have


"A=x^2(y-1);\ B=-x(y^2-1);\ C=y(y-1)"


The roots of the equations "A\u03b1^2+B\u03b1+C= 0" i.e., "x^2(y-1)\alpha-x(y^2-1)\alpha+y(y-1)=0" are given by "λi=±\frac1x".The differential equations for the family of characteristics curves are


"\frac{dy}{dx}\pm \frac1x=0"

Whose solutions are "y+ln\ x=c_1" , and "y-ln\ x=c_2". Choose


"\xi=y+ln\ x,\ \eta=y-ln\ x"

An easy computation shows that

"u_x=u_{\xi}\xi_x+u_{\eta}\eta_x,\\
u_{xx}=u_{\xi \xi}\xi^2_x-2u_{\xi \eta}x^2+u_{\eta \eta}x^2+u_{\xi}-u_{\eta},\\
u_{yy}=u_{\xi \xi}+2u_{\xi \eta}+u_{\eta \eta}"

Substituting these expression in the equation "x^2(y-1)u_{xx}-x(y^2-1)u_{xy}+y(y-1)u_{yy}+(xy)u_x-u_y =0" yields the conical form and the solution to the PDE.